OpenStudy (anonymous):

Evaluate 80 × a^-2 for a = 4.

OpenStudy (anonymous):

its multiple chose answers are 20 5 1/20 1/5

OpenStudy (aravindg):

this is equivalent ti \[80 \times 1/4^2=?\]

OpenStudy (aravindg):

*to

OpenStudy (anonymous):

?

OpenStudy (aravindg):

whats ur doubt?

OpenStudy (anonymous):

im not understanding what you want me to figure out

OpenStudy (aravindg):

a^-2=1/a^2

OpenStudy (aravindg):

we need to evaluate for a=4

OpenStudy (aravindg):

putting a=4 we get above expression then divide and get answer

OpenStudy (anonymous):